OpenStudy (anonymous):

How many natural numbers \(n\) are there such that\[\sqrt{n+\sqrt{n+\sqrt{n+\sqrt{n+...}}}}\]is natuaral too.

hartnn (hartnn):

p^2 - p - n =0

OpenStudy (anonymous):

n=2,6 gives 2 and 3 ,... ??? are there more numbers?

OpenStudy (anonymous):

 then, a^2=n+a